What mass of water will be produced from 9.6g of calcium hydroxide reacting with HCl?
1
Expert's answer
2020-05-08T14:10:22-0400

Ca(OH)2 + 2HCl=CaCl2 + 2H2O

9.6*36/74=‭4,67027027027027‬ g